40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 3645     Entry time: Mon Oct 4 19:48:00 2010
Author: yuta 
Type: Update 
Category: VAC 
Subject: several mirrors installed to ITMX/BS chamber 

(Koji, Kiwamu, Yuta)

Lots of progress for the optics placement in the vacuum chamber.
We are ready to go to the next step: open the access connector between IMC and OMC.
 


Background:

The actual work in the vacuum chamber started.
The first goal of the vent is to get the green beam coming out from the chambers to the PSL table.


What we did:

1. Inspection of the tip-tilt suspension

The alignments of the TTs were inspected. We had five TTs having been sitting on the table in Bob's clean room.
Prior to their installation we checked the alignments of those because they sometimes showed large hysteresis mainly in the pitch directions.

If a TT has the hysteresis, the pitch position doesn't come back to the previous position. This may cause an issue of the interferometer operation.

- SN001, SN003, SN005 looked well aligned and show no hysteresis.

- The alignment of SN002 was not so good (theta ~ 0.004 rad ) compared to those three, but no hysteresis, we think this guy is still acceptable for the installation.

- SN004 had an apparent hysteresis. This guy doesn't come back to the previous place, being applied a touch. We have to fix this at some point.

2. Work in the ITMX chamber

Now all of the optic in this chamber was installed in the approximate place.

- Installed POX/POP steering mirrors.

- TT(SN003) was placed.

- The two steering optics for ITMX OPLEV were placed at the designed positions.

3. Work in the BS chamber

Installed 2 TTs to the BS chamber.
- SR3: TT(SN001)
- PR3: TT(SN005)

After the alignment of the green steering mirrors, we confirmed
the green beam is successfully hitting the west wall of the OMC chamber.

Those TTs are approximately aligned using the weakly reflected green beams.

Next work:

- Open the access connector

- Place another periscope and two steering mirrors for green

- Damping of the suspended optics

- Resurrect MC and its stable lock

- Remove MCT pickoff path

- Align optics in the main path

- Recycled Michelson lock

 

Attachment 1: P1060901.JPG  2.403 MB  Uploaded Wed Oct 6 08:49:30 2010  | Hide | Hide all
P1060901.JPG
Attachment 2: P1060904.JPG  2.375 MB  Uploaded Wed Oct 6 08:50:11 2010  | Hide | Hide all
P1060904.JPG
ELOG V3.1.3-